OpenStudy (brinazarski):

The expression cos(A-B) - cos(A+B) is equal to 1) -2sinAsinB 2) -2cosB 3) 2cosAcosB 4) 2sinAsinB So far, all I've gotten is cosxcosy + sinxsiny - cosxcosy - sinxsiny I don't know where to go from there. I'm assuming that's zero, and I don't know which of the others equal 0... they all look the same.

OpenStudy (sasogeek):

well here's what u did wrong, -cos(A+B)=-(cosAcosB-sinAsinB)=-cosAcosB+sinAsinB ... try reworking with this...

OpenStudy (brinazarski):

It's not -cos(A+B)... It's just subtracting +cos(A+B), isn't it? So that shouldn't change anything...

OpenStudy (sasogeek):

well u see, if you have a-(b-c), it is the same as a-(+b-c) but when u try to expand this, it becomes a-b+c, that's because two negatives will make a positive :)

OpenStudy (brinazarski):

Sooo... when you subtract cos(x + y), it turns into adding cos(x-y)....?

OpenStudy (sasogeek):

cos(A+B) will give you cosAcosB-sinAsinB , i agree to this, let's substitute this into the expression cos(A-B) - \(cos(A+B)\)=cosAcosB+sinAsinB - \( (cosAcosB-sinAsinB) \) cosAcosB+sinAsinB-cosAcosB+sinAsinB remember, two negatives make a positive you'll have the cosAcosB-cosAcosB=0 and sinAsinB+sinAsinB=2sinAsinB

OpenStudy (sasogeek):

do u get it though?

OpenStudy (brinazarski):

Sorry, someone came over here :/

OpenStudy (brinazarski):

I think I get it, it's kind of weird though...

OpenStudy (sasogeek):

-(a-b) is not the same as -a-b, here's why, -(a-b) is the same as -1(a-b), if you try and use the distribution property, then you'll have (-1a)+(-1(-b)) which will become -a+b because -1(-b)=b two negatives make a positive... so in a nut shell, -(-x)=x try and apply that :) it will make sense :)